OpenStudy (anonymous):

A question about using the Gaussian method of Elimination to find a Determinant. When subtracting/adding rows in the following form: R2 → R2 - 2*R1 The determinant's value remains. However when I have to multiply the rows I'm changing, the value of the determinant changes completely. Why is that so? Examples below:

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 4 & 3 & 2 \\ 2 & 1 & 4 & 3 \\ 3 & 2 & 1 & 4 \\ 4 & 3 & 2 & 1\end{matrix}\right]\] det = -120 \[R_2 \rightarrow R_2 -2R_1\] \[R_3 \rightarrow R_3 -3R_1\] \[R_4 \rightarrow R_4 -4R_1\] \[\left[\begin{matrix}1 & 4 & 3 & 2 \\ 0 & -7 & -2 & -1 \\ 0 & -10 & -8 & -2 \\ 0 & -13 & -11 & -6\end{matrix}\right]\] det = -120 --------------------------------------------------- \[\left[\begin{matrix}2 & -3 & 0 & 1 \\ 1 & 5 & -1 & 2 \\ -1 & 0 & 4 & -2 \\ 3 & 2 & 7 & 6\end{matrix}\right]\] det = 267 \[R_2 \rightarrow 2R_2 - R_1\] \[R_3 \rightarrow 2R_3 + R_1\] \[R_4 \rightarrow 2R_4 - 3R_1\] \[\left[\begin{matrix}2 & -3 & 0 & 1 \\ 0 & 13 & -2 & 3 \\ 0 & -3 & 8 & -3 \\ 0 & 13 & 14 & 9\end{matrix}\right]\] det = 2136

OpenStudy (anonymous):

I triple checked the row elimination calculations, but the inconsistencies always seem to return when I'm multiplying the row I'm changing by a scalar.

OpenStudy (solomonzelman):

Actually, when you multiply a row by something (i.e. when you replace \(R_j\) by \(c\cdot R_j\)), you should divide the determinant by \(c\), and when you switch row, multiply the determinant by -1. When you add rows, however, you don't change the determinant.

OpenStudy (solomonzelman):

Your steps should look like \(\color{#000000}{ \displaystyle a\cdot R_j +R_i\rightarrow R_i }\) Don't do, \(\color{#000000}{ \displaystyle a\cdot R_j +bR_i\rightarrow R_i }\) because you will get confused and forget to multiply the determinant by scalar.